OpenStudy (anonymous):

For the function f(x) = (8-2x)2 ,find f-1 . Determine whether f-1 is a function.

OpenStudy (anonymous):

@Catlover5925

OpenStudy (m3ga-bit3):

@ganeshie8

OpenStudy (m3ga-bit3):

i suck at math sry:(

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Catlover5925 @...Dimps...

OpenStudy (anonymous):

solve \[x=(8-2y)^2\] for \(y\) first step is \[8-2y=\pm x\] so it will definitely not be a function

OpenStudy (anonymous):

ok lol i understand it some what i just have trouble finding the answer

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

after you get to \[8-2y=\pm x\] it takes a couple steps to solve for \(y\) but the \(\pm\) tells you it is not a function to solve for \(y\) could be \[8-2y=\pm x\\ -2y=\pm x-8\\ y=\pm\frac{x}{2}+4\]

OpenStudy (anonymous):

@Directrix

OpenStudy (m3ga-bit3):

@Abhisar

OpenStudy (anonymous):

there areno y's in any of my answers

OpenStudy (m3ga-bit3):

y they even have letters in math any way?!?!?

OpenStudy (anonymous):

idkbut high school math will fry ya brain